# Pretraining a Neural Operator in Lower Dimensions

### v1

@@ 4.1 Datasets and objectives
The general format and domains of the PDEs in this work are shown in equation 6, where $N$ is a linear or nonlinear function, $\mathbf{c}$ is the vector of the PDE coefficients, $u$ is the quantity of interest, and $t,\mathbf{x}$ are the time and space variables, respectively. Each sample of a dataset consists of the evolution of the system in the specified time domain sampled every $\Delta t=0.05$, resulting in 21 snapshots and 20 input-output pairs on which the model will be trained. The spatial resolutions for the 1D and 2D datasets are set to $1024$ and $64^{2}$, respectively. Different samples of each dataset are governed by the same equation and coefficients, and differ only in the initial condition defined by equation 7 where $\mathbf{A}_{i},n_{i},\phi_{i}$ are random variables drawn from uniform distributions, similar to the multi-dimensional datasets in PDEBench [49]. We assume periodic boundary conditions in all cases.
@@
For each dimensionality and each value of $\mathbf{c}$, 10000 samples were generated, 2000 of which are held for validation. We use the entire 1D training dataset in the pretraining stage, but use a limited subset of the 2D training data for the downstream task. Since the characteristics of a typical scenario for using pretraining models are the high collection cost and scarcity of data, we set up the experiments to resemble such scenarios.
For the downstream task, we compare a model trained from random initialization and 8 pretrained models with fine-tuning configurations shown in table 1. The objective of both training and validation is the relative $L_{2}$ norm (also known as nRMSE [49]) of the prediction error shown in equation 8 where $\hat{y},y$ are the model output and target value for the quantity of interest, respectively. In our application, $y$ is the state of the system at the next time-step. We train the model on the next-step prediction loss and evaluate the model using the same metric on the validation dataset, as well as the average loss over a 5-step autoregressive rollout similar to what was done in PDEBench [49]. Finally, we run each experiment with 3 different random seeds and report the average result to take into account the randomness of the network initialization, as well as the randomly selected subset of the 2D training dataset. Now, we will look at the specifics of each equation, the result for each dataset, and discuss the effect of the fine-tuning strategy, the number of training samples, and the PDE coefficient based on the observed results. We will continue the discussion here using concise plots. Please refer to A for detailed numerical tables on all the curated results.
@@
Each training stage consists of 5000 iterations of the AdamW optimizer with an initial learning rate of 0.001. The learning rate is multiplied by 0.2 when the loss reaches a plateau and does not improve for more than 100 iterations. Our choice of architecture for FFNO has 4 hidden layers with a latent dimension of 128 and 16 Fourier modes in each axis.
@@ 4.2 Diffusion equation
The results for the diffusion datasets are shown in figure 2. The x-axis represents the number of downstream training samples, and the y-axis represents the average loss on the 2000 validation samples. We exclude the instances trained with only 1 or 2 training samples due to their unacceptable amount of error. It appears that the fine-tuning configurations C1, C2, and C8 perform as well or better than the randomly initialized model C0, and the remaining configurations fail to achieve the same error as C0. Focusing on C1 (all parameters trainable) and C0, the first trivial indicator of success is that C1 outperforms C0 and reduces the error up to 80% in the low-data regime. The second interesting observation is the different trend of the error with respect to the number of training parameters. The error of C0 starts from a very high value in the left side with very few samples, and improves significantly with the increase of the training dataset size. However, the performance of C1 even with very few samples on the left is already comparable to the performance on the right end of the plot. An interesting observation is that C1 trained on 8 samples outperforms C0 trained on 1024 samples for $\nu=0.004$. There we can also see more clearly that the advantage of pretraining is amplified over autoregressive rollout. For exact values, refer to A.
@@
Focusing on the fine-tuning configurations, C1, C2, and C8 are the models that perform best. The explanation is trivial for C1, since it is totally free to be tuned on the downstream data. However, it is interesting that C8 with only the last layer left trainable outperforms C2 (all Fourier layers trainable) despite having fewer trainable parameters. Each Fourier and feedforward layer have about 524K and 60K parameters, respectively. This means that C8 with less than 600K parameters performs as well as or better than C2 with more than 2 million trainable parameters. This can be indicative of a hierarchical order of the learning mechanism of the model for the diffusion equation. In the extreme low-data regime (less than 10), tuning configurations with less trainable parameters appear to have the most advantage according to the tables in A. This makes sense since having fewer trainable parameters reduces their risk of overfitting. However, they fall behind when more training samples are available.
@@ 4.3 Advection equation
Similarly to the diffusion dataset, six advection datasets were provided for our experiments. The 1D datasets are provided by PDEBench [49], and we generated the 2D datasets using the exact solution function of $u(x,y,t)=u_{0}(x-\beta t,y-\beta t)$.
@@ 5 Conclusion
This work proposes a novel strategy to reduce the cost of data collection, training, and error of a neural operator in multidimensional PDEs. For a neural operator with a factorized architecture like FFNO, we can pretrain a model in lower dimensions (PreLowD) and transfer the weights to a multidimensional model. We showed that this strategy is successful for the 2D diffusion equation, but not for the advection equation. When successful, the PreLowDed model outperformed the randomly initialized model by up to 70% in low-data regimes and a slow-changing system. When the system has a faster rate of change or the model is used autoregressively, the gain of the PreLowDed model is even more significant, achieving an error reduction of 50% over a 5-step rollout with 1024 training samples.
We explored several fine-tuning strategies to find the best way to freeze or tune the transferred weights from a PreLowDed model. It seems that with sufficient training, more trainable parameters will result in a lower error. However, tuning a smaller subset of the parameters may achieve a lower error when very few training samples are available. For the advection equation, the only important factor seemed to be the sheer number of tunable parameters. However, tuning the final full layer for the diffusion datasets was among the best choices, indicating a possibly hierarchical feature extraction of the model.
In future work, this pretraining strategy may be utilized with other neural operators and PDE solvers that have the necessary properties as well. However, it may not be straightforward to find or define a version of the system of interest in a different dimensionality. Moreover, the selective fine-tuning strategy can be applied to any neural operator and can provide insight into the generalizability and interpretability of such models. For example, tuning different parts of the model while varying a certain coefficient or term for the PDE of the downstream task can help discover the correspondence of the neural operator’s components and different modalities and terms of the PDE. Finally, an important potential venue for research is the study of different possible compositions of low-dimensional models and factorization of high-dimensional models. The ultimate goal of this direction is to improve the prediction accuracy and efficiency in more complex or 3D systems, and this work was just a humble starting step towards that end.
Appendix A All results

### v2

@@ 5.1 Datasets and objectives
The general format and domains of the PDEs in this work are shown in equation 6, where $N$ is a linear or nonlinear function, $\mathbf{c}$ is the vector of the PDE coefficients, $u$ is the quantity of interest and $t,\mathbf{x}$ are the time and space variables, respectively. Each sample of a dataset consists of the evolution of the system in the specified time domain sampled every $\Delta t=0.05$, resulting in 21 snapshots and 20 input-output pairs on which the model will be trained. The spatial resolutions for the 1D and 2D datasets are set to $1024$ and $64^{2}$, respectively. Different samples of each dataset are governed by the same equation and coefficients, and differ only in the initial condition defined by equation 7 where $\mathbf{A}_{i},n_{i},\phi_{i}$ are random variables drawn from uniform distributions, similar to the multi-dimensional datasets in PDEBench (Takamoto et al., 2022). We assume periodic boundary conditions in all cases.
@@
For each dimensionality and each value of $\mathbf{c}$, 10000 samples were generated, 2000 of which are held for validation. Since the characteristics of a typical scenario for pretraining a model are the high collection cost and scarcity of data, we set up the experiments to resemble such scenarios. For each case (a particular PDE with a particular value for the coefficient), we first pretrain the model on the relevant 1D pretraining dataset. Then, we fine-tune the pretrained model with different amounts of available downstream training data. That is to find out the relationship between downstream dataset size and the gain of pretraining compared to training from scratch. To clarify, the experiment for each PDE and each value for the coefficient(s) of the PDE are separate and independent of each other. For example, if the downstream task is to learn 2D diffusion with a diffusion coefficient of $\nu=0.002$, the pretraining dataset consists only of samples from 1D diffusion equation with the same diffusion coefficient of $\nu=0.002$.
For the downstream task, we compare a model trained from random initialization and 8 pretrained models with fine-tuning configurations shown in table 1. The objective of both training and validation is the relative $L_{2}$ norm, also known as nRMSE (Takamoto et al., 2022), of the prediction error shown in equation 8 where $\hat{y},y$ are the model output and target value for the quantity of interest, respectively. In our application, $y$ is the state of the system at the next time-step. We train the model on the next-step prediction loss and evaluate the model using the same metric on the validation dataset, as well as the average loss over a 5-step autoregressive rollout similar to what was done in PDEBench (Takamoto et al., 2022). Finally, we run each experiment with 3 different random seeds and report the average result to take into account the randomness of the network initialization, as well as the randomly selected subset of the 2D training dataset. Now, we will look at the specifics of each equation, the result for each dataset, and discuss the effect of the fine-tuning strategy, the number of training samples, and the PDE coefficient based on the observed results. We will continue the discussion here using concise plots. Please refer to Appendix A for detailed numerical tables on all the curated results.
@@
Each training stage consists of 5000 iterations of the AdamW optimizer with an initial learning rate of 0.001. The learning rate is multiplied by 0.2 when the loss reaches a plateau and does not improve for more than 100 iterations. Our choice of architecture for FFNO has 4 hidden layers with a latent dimension of 128 and 16 Fourier modes in each axis.On our GeForce RTX 2080 Ti Nvidia GPUs, with a batch size of 64 and 5000 optimization iterations for each stage, the pretraining takes about 3.35 minutes while the fine-tuning or training for the 2D task takes about 21 minutes. This is an example of how the training cost differs for a 2D PDE compared to 1D. This is despite the resolution in 2D being 64 while the 1D data has a resolution of 1024.
@@ 5.2 Diffusion equation
The results for the diffusion datasets are shown in figure 2. The x-axis represents the number of downstream training samples, and the y-axis represents the average loss on the 2000 validation samples. We exclude the instances trained with only 1 or 2 training samples due to their unacceptable amount of error. It appears that the fine-tuning configurations C1, C2, and C8 perform as well or better than the randomly initialized model C0, and the remaining configurations fail to achieve the same error as C0. Focusing on C1 (all parameters trainable) and C0, the first trivial indicator of success is that C1 outperforms C0 and reduces the error up to 80% in the low-data regime. The second interesting observation is the different trend of the error with respect to the number of training parameters. The error of C0 starts from a very high value in the left side with very few samples, and improves significantly with the increase of the training dataset size. However, the performance of C1 even with very few samples on the left is already comparable to the performance on the right end of the plot. An interesting observation is that C1 trained on 8 samples outperforms C0 trained on 1024 samples for $\nu=0.004$. There we can also see more clearly that the advantage of pretraining is amplified over autoregressive rollout. For exact values, refer to Appendix A.
@@
Focusing on the fine-tuning configurations, C1, C2, and C8 are the models that perform best. The explanation is trivial for C1, since it is totally free to be tuned on the downstream data. However, it is interesting that C8 with only the last layer left trainable outperforms C2 (all Fourier layers trainable) despite having fewer trainable parameters. Each Fourier and feedforward layer have about 524K and 60K parameters, respectively. This means that C8 with less than 600K trainable parameters performs as well as or better than C2 with more than 2 million trainable parameters. This can be indicative of a hierarchical order of the learning mechanism of the model for the diffusion equation. In the extreme low-data regime (less than 10), tuning configurations with less trainable parameters appear to have the most advantage according to the tables in Appendix A. This makes sense since having fewer trainable parameters reduces the risk of overfitting. However, they fall behind when more training samples are available.
@@ 5.3 Advection equation
Similarly to the diffusion dataset, six advection datasets were provided for our experiments. The 1D datasets are provided by PDEBench (Takamoto et al., 2022), and we generated the 2D datasets using the exact solution function of $u(x,y,t)=u_{0}(x-\beta t,y-\beta t)$.
@@ 6 Conclusion
This work proposes a novel strategy to reduce the cost of data collection, training, and error of a neural operator in multidimensional PDEs. For a neural operator with a factorized architecture like FFNO, we can pretrain a model in lower dimensions (PreLowD) and transfer the weights to a multidimensional model. We showed that this strategy is successful for the 2D diffusion equation, but not for the advection equation. When successful, the PreLowDed model outperformed the randomly initialized model by up to 80% in low-data regimes and a slow-changing system. When the system has a faster rate of change or the model is used autoregressively, the gain of the PreLowDed model is also very significant, achieving an error reduction of 50% over a 5-step rollout even with 1024 training samples.
We explored several fine-tuning strategies to find the best way to freeze or tune the transferred weights from a PreLowDed model. It seems that with sufficient training, more trainable parameters will result in a lower error. However, tuning a smaller subset of the parameters may achieve a lower error when very few training samples are available. For the advection equation, the only important factor seemed to be the sheer number of tunable parameters. However, tuning the final full layer for the diffusion datasets was among the best choices, indicating a possibly hierarchical feature extraction of the model. Our general recommendation is to fine-tune the final layer in case of extreme scarcity of training data in the downstream task, or to fine-tune the whole model with all parameters being trainable.
In future work, this pretraining strategy may be utilized with other neural operators and PDE solvers that have the necessary properties as well. However, it may not be straightforward to find or define a similar physical system with fewer number of dimensions for pretraining. Moreover, the selective fine-tuning strategy can be applied to any neural operator and can provide insight into the generalizability and interpretability of such models. For example, tuning different parts of the model while varying a certain coefficient or term for the PDE of the downstream task can help discover the correspondence of the neural operator’s components and different modalities and terms of the PDE. Finally, an important potential venue for research is the study of different possible compositions of low-dimensional models and factorization of high-dimensional models. The ultimate goal of this direction is to improve the prediction accuracy and efficiency in more complex or 3D systems, and this work was just a humble starting step towards that end.
Appendix A Appendix